OpenStudy (anonymous):

Calculate the molarity of 29.0g of ethanol (C2H5OH) in 545 mL of solution? I'm not sure what steps to use

OpenStudy (callisto):

no of mole of C2H5OH = 29.0 / (12*2+1.0*6+16.0) = 0.630434782mole molarity = no of mole/volume = 0.630434782 /(500/1000) = 1.26moldm^-3 = 1.26M

OpenStudy (anonymous):

First step is to work out how many moles you have. 1 mole is a formula weight expressed in grams, so as Callisto correctly points out it's 2 times 12 for the carbons plus 6 times 1 for the hydrogens and 1 times 16 for the oxygen. This adds up to 46. If you had 46 grams in a litre of solution then it would be a 1M solution but you have less than that: 0.63 moles. However the volume is less than 1 litre so the concentration will be 0.63 times 1000 divided by 545 which I make to be 1.157M. I hope this working out helps you understand how it's done. (PS The weights of the individual atoms are on the periodic table).

OpenStudy (anonymous):

I see now! Thank you both so much! I really appreciate it! And yes, this did help.

OpenStudy (callisto):

sorry, i sub the wrong value :(